The left and right page numbers of an open book are two consecutive integers whose sum is 519.Find these page numbers.
1
Expert's answer
2019-08-30T08:48:29-0400

Solution

Let x be the left number, then x+1 the right number. According to the condition,the sum of two numbers is 519, hence


"x+(x+1)=519."

Solve this equation


"x+x+1=519;"


"2x+1=519;"


"2x=518;"


"x=259;"


"x+1=260"

Answer:

The number of the left page is 259 and the number of the right page is 260.